OpenStudy (twistrose):

Juan deposited $200 in a savings account earning 3% interest over 2 years. What was the total amount of interest earned at the end of 2 years? A. $12 B. $24 C. $120 D. $240

OpenStudy (knov):

200/100=x/3 100x=600 x=6 over 2 years which means x=2*6 and it's 12 dollars.

OpenStudy (twistrose):

thanks

OpenStudy (knov):

no problem